# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v2

@@ Abstract
There have been many parallel streams of research studying order isomorphisms of some specific sets $\mathcal{G}$ of functions from a set $\mathscr{X}$ to {$\mathbb{R}\cup\{\pm\infty\}$, such as the sets of convex or Lipschitz functions. We propose in this article a unified approach inspired by $c$-convex functions. Our results are obtained highlighting the role of inf and sup-irreducible elements of $\mathcal{G}$ and the usefulness of characterizing them, to subsequently derive the structure of order isomorphisms, and in particular of those commuting with the addition of scalars. We show that in many cases all these isomorphisms $J:\mathcal{G}\to\mathcal{G}$ are of the form $Jf=g+f\circ\phi$ for a translation $g:\mathscr{X}\to\mathbb{R}$ and a bijective reparametrization $\phi:\mathscr{X}\to\mathscr{X}$. Given a reference anti-isomorphism, this characterization then allows to recover all the other anti-isomorphisms. We apply our theory to the sets of $c$-convex functions on compact Hausdorff spaces, to the set of lower semicontinuous (convex) functions on a Hausdorff topological vector space and to 1-Lipschitz functions of complete metric spaces. The latter application is obtained using properties of the horoboundary of a metric space.
@@
Let $\mathscr{X}$ and $\mathscr{Y}$ be nonempty sets. Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a subset of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. ¯$\overline{\mathbb{R}}^{\mathscr{Y}}$), with both $\mathscr{F}$ and $\mathcal{G}$ being proper and separating, stable by arbitrary suprema and addition of scalars. Let $J$ be a (max,+)-isomorphism from $\mathscr{F}$ onto $\mathcal{G}$ and define $e_{x}=\sup_{u\in\mathcal{G}}u(\cdot)-u(x)$ and $e^{\prime}_{y}=\sup_{v\in\mathscr{F}}v(\cdot)-v(y)$. Then the following statements are equivalent:
@@
The precise statement is to be found on 21, proper and separating corresponding to 1, and the functions $e_{x}$ being introduced in (18). The other properties of $g$ and $\phi$ depend obviously on the specific $\mathscr{F}$ and $\mathcal{G}$ studied, some characterizations obtained are regrouped in Table 1. The particularity of our approach is that we encompass all these functions spaces within a single abstract formalism of sup-stable spaces. 21 establishing (1) is only a posteriori adapted to characterize the $g$ and $\phi$ for each specific space.
@@
1-Lipschitz functions constant isometry 34
@@
42
Concerning order reversing isomorphisms $T$ between sets $\mathscr{F}$ and $\mathcal{G}$ commuting with the addition of scalars, we characterize their existence through kernels $b$ in 10. Owing to the immediate 11, the uniqueness of $T$ up to basic transformations is instead a consequence of results like 1.
On $c$-convex functions over a compact set, we show in 27 that the existence of a (max,+)-isomorphism between $b$-convex and $c$-convex functions is equivalent to a “conjugacy” formula (25) satisfied by the kernels. Concerning 1-Lipschitz functions, the results rely on the compactification of metric spaces by horofunctions, and in particular on properties of Busemann functions (36), we also allow for weak distances, i.e. potentially assymetric and with negative values.
Similarly to the proof of the Banach-Stone theorem, see e.g. the lecture notes [GJ02], the key idea in our proofs is to advocate for inf and sup-irreducibility as a good notion of “extremality” (4) and to show that these “extremal” points are preserved by order isomorphisms (13) and generate the whole space, in a Krein–Milman fashion (19 and 25).
Beyond (max,+)-isomorphisms, we study the order isomorphisms of sets containing the indicator functions of points, such as the set of lower semicontinuous functions on Hausdorff topological spaces (30) and its subset of convex functions (42). In these cases, indicator functions play a major role in the characterization. We in particular obtain an extension of the theorems of [AAM09, IRS15] in the generality of locally convex Hausdorff spaces.
@@
Some of our results have precursors when $\mathscr{X}$ is finite, in the setting of tropical geometry – the (max,+) structure considered here being the main example of tropical algebraic structure. In particular, the notion of irreducibility was used in [BSS07, GK07] when studying the max-plus analogue of the representation of convex sets in terms of extreme points or extreme rays. Concerning infinite sets, the literature can be regrouped according to the functions spaces studied:
First, on convex functions, our focus meets that of the research program set by S. Artstein-Avidan and V. Milman on the characterization of some operations on subsets of convex functions. In their seminal paper [AAM09], they have shown that, for $\mathcal{G}$ the space of convex lower semicontinuous functions over $\mathbb{R}^{d}$, the only possible anti-involutions are compositions of the Fenchel transform with affine transformations. This result was further extended to Banach spaces by [IRS15], followed by [CL21] who characterized for which Banach spaces the Fenchel transform is an anti-involution. We refer to [IRS15, AASW23] for a list of other developments on $s$-concave functions or convex bodies. Underlying all of these works on convex functions is a fundamental result of affine geometry that, in dimensions larger than 2, transformations preserving straight lines are affine. The notable difference of our approach is that we consider spaces beyond convex functions on vector spaces, so in general we do not have linear structures.
@@
Finally, an important application of ours is to the set of Lipschitz functions on a complete metric space with Lipschitz constant at most 1. Most of the interest for spaces of Lipschitz functions has been for their Banach structure, but they also have a rich lattice structure as advocated by [Wea94]. This induces quite some polisemy on the notion of isomorphism. For instance [CCD19] studies Banach space isomorphisms between spaces of Lipschitz functions. Weaver leveraged instead the inf/sup-lattice properties of the vector spaces of Lipschitz functions to characterize linear Banach space isomorphisms preserving Riesz-norms [Wea94]. Banach–Stone-like theorems for Lipschitz functions have also been shown for linear isomorphisms such as [JVCVV09, Theorem 4.1] and nonlinear [CSCS11, Theorem 1] and references within. Finally Leung et al. for linear, in [LL13], and nonlinear order isomorphisms, in [LT16], designed an abstract framework to describe order isomorphisms on “near vector lattices”. However all the aforementioned authors rely extensively on vector subspaces of the space of continuous functions, while we consider sup-stable sets of functions. The closest research to our work was done by [CSJ17, DJVM20] for $1$-Lipschitz functions on metric spaces with non-symmetric distance with the assumption that the isomorphisms also preserve convex combinations. Nevertheless, none of the authors uses the notions of inf or sup-irreducible functions, whereas these are the backbone of our analysis.
There is a separate direction of research to characterize cones for which every order isomorphim is affine (in $f$). We refer to [LvGvI20, Wal19] for some technical conditions ensuring this. They unfortunately are not sufficient on spaces like that of continuous functions if not paired with homogeneity assumptions on $J$, in line with the work of [Sch79].
@@
In Section 1 we introduce the notions of irreducibility and of isomorphisms that we consider. Section 2 pertains to our main results on order isomorphisms, in particular when commuting with constants, with Section 2.3 devoted to $c$-convex functions. In Section 3, we detail the consequences of these results for continuous and lower semicontinuous functions (Section 3.1), for 1-Lipschitz functions (Section 3.2), and finally for convex lower semicontinuous functions (Section 3.3).
@@ 1 Preliminaries
Notation: The extended real line is denoted by $\overline{\mathbb{R}}=[-\infty,+\infty]$, we also use the notations $\mathbb{R}_{\top}=(-\infty,+\infty]$, $\mathbb{R}_{\bot}=[-\infty,+\infty)$ and $\mathbb{R}_{>0}=(0,+\infty)$. Given any sets $\mathscr{X}$ and $E$, we denote by $E^{\mathscr{X}}$ the set of functions $f$ from $\mathscr{X}$ to $E$ and $\operatorname*{Id}_{\mathscr{X}}:\mathscr{X}\to\mathscr{X}$ the identity map. When $\mathscr{X}$ is equipped with a topology, a function $f:\mathscr{X}\rightarrow\overline{\mathbb{R}}$ is said to be lower semicontinuous (l.s.c.) if its epigraph is a closed subset of $\mathscr{X}\times\overline{\mathbb{R}}$. Its domain is defined by $\operatorname*{Dom}(f)=\{x\in\mathscr{X}\,|\,f(x)<+\infty\}$. A function $f$ is said to be proper if its domain is non-empty and if it does not take the value $\infty$- ∞. For $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ and some $x\in\mathscr{X}$, if there exists a function $g\in\mathcal{G}$ such that $g(x)\in\mathbb{R}$, we say that $\mathcal{G}$ is proper at $x$, and more simply that $\mathcal{G}$ is proper if it is at all $x\in\mathscr{X}$. When $\mathscr{X}$ and $\mathscr{Y}$ are topological vector spaces, the set $\mathscr{X}^{*}$ denotes the dual space of $\mathscr{X}$ of continuous linear forms from $\mathscr{X}$ to $\mathbb{R}$ and $\mathcal{L}(\mathscr{X},\mathscr{Y})$ refers to the space of continuous linear operators from $\mathscr{X}$ to $\mathscr{Y}$. We write the duality product over $\mathscr{X}^{*}\times\mathscr{X}$ as $\langle\cdot,\cdot\rangle:\mathscr{X}^{*}\times\mathscr{X}\to\mathbb{R}$.
@@
We say that a subset $\mathcal{G}$ of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ is sup-stable (resp. finitely sup-stable) if it is stable under arbitrary sups (resp. pairwise sups), and that it is a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ if it is stable under arbitrary sups and addition of scalars. For $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$, we define its sup and inf-closures as
$\overline{\mathcal{G}}^{\sup}$ :=$:=\{\sup_{\alpha\in\mathcal{A}}h_{\alpha}\,|\,\mathcal{A}\text{ an index set},\,\{h_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{G}\}$
$\overline{\mathcal{G}}^{\inf}$ :=$:=\{\inf_{\alpha\in\mathcal{A}}h_{\alpha}\,|\,\mathcal{A}\text{ an index set},\,\{h_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{G}\}$
@@
Relative infima underline that infima or suprema can be done either w.r.t. the absolute (entry-wise) supremum of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ or its relative counterpart in $\mathcal{G}$.
@@
Let $\mathcal{G}$ be a subset of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$. We say that a point $f\in\mathcal{G}$ is sup-irreducible if, for all $g,h\in\mathcal{G}$,
@@
Let $\mathcal{G}$ be a sup-stable subset of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$. For all $\tilde{g},\tilde{h}\in\overline{\mathbb{R}}^{\mathscr{X}}$, if $f=\inf(\tilde{g},\tilde{h})$ with $f\in\mathcal{G}$, then $f=\inf\nolimits^{\mathcal{G}}(\inf\nolimits^{\mathcal{G}}\tilde{g},\inf\nolimits^{\mathcal{G}}\tilde{h})$. The $\mathcal{G}$-relatively-inf-irreducible functions are thus also inf-irreducible.
@@
Take $f=\inf(\tilde{g},\tilde{h})$ with $f\in\mathcal{G}$. Since $f\leq\tilde{g}$, we have that $f\leq\inf\nolimits^{\mathcal{G}}(\tilde{g})$ by definition of $\inf\nolimits^{\mathcal{G}}$ in (2). The same holds for $\tilde{h}$, hence $f\leq\tilde{f}:=\inf(\inf\nolimits^{\mathcal{G}}\tilde{g},\inf\nolimits^{\mathcal{G}}\tilde{h})$ and $f\leq\inf\nolimits^{\mathcal{G}}\tilde{f}\leq\tilde{f}$. Moreover $\tilde{f}\leq\tilde{g}$ and $\tilde{f}\leq\tilde{h}$, so $\tilde{f}\leq f=\inf(\tilde{g},\tilde{h})$ and $f=\tilde{f}$. If $\tilde{g},\tilde{h}\in\mathcal{G}$ and $f$ is $\mathcal{G}$-relatively-inf-irreducible, then w.l.o.g $f=\inf\nolimits^{\mathcal{G}}\tilde{g}=\tilde{g}$ so $f$ is inf-irreducible. ∎
@@
$\delta^{\bot}_{x}\left(y\right):=\left\{\begin{array}[]{ll}0&\text{ if }y=x,\\ -\infty&\text{ otherwise,}\end{array}\right.\quad\delta^{\top}_{x}\left(y\right):=\left\{\begin{array}[]{ll}0&\text{ if }y=x,\\ +\infty&\text{ otherwise.}\end{array}\right.$ (6)
@@
    a (max,+)-isomorphism if it is a max-isomorphism and if we have $J(f+\lambda)=Jf+\lambda$ for $\lambda\in\mathbb{R}$, assuming that $\mathcal{G}$ and $\mathscr{F}$ are complete subspaces of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$;
@@
    an anti-involution if $J:\mathcal{G}\rightarrow\mathcal{G}$, $JJ=\operatorname*{Id}_{\mathcal{G}}$ and $J$ is order reversing.
@@
Let ¯$\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ be sup-stable sets. Then the order isomorphisms $J:\mathscr{F}\rightarrow\mathcal{G}$ commute with an arbitrary suprema, and thus coincide with max-isomorphisms. Moreover, if $J:\mathscr{F}\rightarrow\mathcal{G}$ is a max-isomorphism, then so is $J^{-1}:\mathcal{G}\rightarrow\mathscr{F}$. Anti-isomorphisms $\bar{F}:\mathscr{F}\to\mathcal{G}$ send relative infima of $\mathscr{F}$ onto suprema of $\mathcal{G}$, i.e. $\bar{F}(\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha})=\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha})$ for any $\{f_{\alpha}\}_{\alpha\in\mathscr{A}}\subset\mathscr{F}$.
@@
So $\mathscr{F}(\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha})=\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha})$.
@@
A map $B:\overline{\mathbb{R}}^{\mathscr{Y}}\rightarrow\overline{\mathbb{R}}^{\mathscr{X}}$ is said to be ¯$\overline{\mathbb{R}}_{\max}$-sesquilinear if $B(\inf\{f_{i}\}_{i\in I})=\sup\{Bf_{i}\}_{i\in I}$ and $B(f+\lambda)=Bf-\lambda$ (with +$\infty$+ ∞ absorbing on the l.h.s. and $\infty$- ∞ absorbing on the r.h.s.), for any finite index set $I$ and $\lambda\in\overline{\mathbb{R}}$; we say in addition that $B$ is continuous if $B(\inf\{f_{i}\}_{i\in I})=\sup\{Bf_{i}\}_{i\in I}$ holds even for infinite families.
@@
A map $\bar{B}:\overline{\mathbb{R}}^{\mathscr{Y}}\rightarrow\overline{\mathbb{R}}^{\mathscr{X}}$ is ¯$\overline{\mathbb{R}}_{\max}$-sesquilinear and continuous if and only if there exists a kernel $b:\mathscr{X}\times\mathscr{Y}\rightarrow\overline{\mathbb{R}}$ such that $\bar{B}f(x)=\sup_{y\in\mathscr{Y}}b(x,y)-f(y)$ for all $f\in\overline{\mathbb{R}}^{\mathscr{Y}}$ (with $\infty$- ∞ absorbing). Moreover in this case $b$ is uniquely determined by $\bar{B}$ as $b(\cdot,y)=\bar{B}\delta^{\top}_{y}$.
Hence the relation between ¯$\overline{\mathbb{R}}_{\max}$-sesquilinear and continuous maps, a.k.a. (Fenchel-Moreau) conjugations, and kernels is one-to-one, so it is equivalent to study a map or its kernel. Moreover 9 allows to write
@@
where $b:\mathscr{X}\times\mathscr{Y}\rightarrow\overline{\mathbb{R}}$ is the kernel associated with $\bar{B}$. Notice also that any complete subspace $\mathcal{G}$ of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ is a range indexed by $\mathcal{G}$ itself, since
$\mathcal{G}=\operatorname*{Rg}(B_{\mathcal{G}})=\{\sup_{g\in\mathcal{G}}a_{g}+b_{\mathcal{G}}(\cdot,g)\,|\,a_{g}\in\mathbb{R}_{\bot}\}\text{ with }b_{\mathcal{G}}(\cdot,g)=g(\cdot).$
@@
with the convention that $\infty$- ∞ is absorbing. Interesting examples of sets $\mathcal{G}=\operatorname*{Rg}(B)$ are that of lower semicontinuous $C$-semiconvex functions for $b(x,y)=-\frac{C}{2}\|x-y\|^{2}$ for some $C>0$, or of 1-Lipschitz functions for $b(x,y)=-d(x,y)$. We use a terminology rooted in abstract convexity theory, however these functions are nothing else than the $c$-convex functions in optimal transport theory, taking $c(x,y)=-b(x,y)$. The key relation is that $\bar{B}=\bar{B}\bar{B}^{\circ}\bar{B}$, see e.g.  [AGK05, p.3,Theorem 2.1, Example 2.8]. We provide now a characterization relating ranges and anti-isomorphisms.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. ¯$\overline{\mathbb{R}}^{\mathscr{Y}}$). Then the following statements are equivalent:

### v3

@@ Abstract
There have been many parallel streams of research studying order isomorphisms of some specific sets $\mathcal{G}$ of functions from a set $\mathscr{X}$ to $\mathbb{R}\cup\{\pm\infty\}$, such as the sets of convex or Lipschitz functions. We develop in this article a unified approach inspired by $c$-convex functions. Our results are obtained highlighting the role of inf and sup-irreducible elements of $\mathcal{G}$ and the usefulness of characterizing them, to subsequently derive the structure of order isomorphisms, and in particular of those commuting with the addition of scalars. We show that in many cases all these isomorphisms $J:\mathcal{G}\to\mathcal{G}$ are of the form $Jf=g+f\circ\phi$ for a translation $g:\mathscr{X}\to\mathbb{R}$ and a bijective reparametrization $\phi:\mathscr{X}\to\mathscr{X}$. Given a reference anti-isomorphism, this characterization then allows to recover all the other anti-isomorphisms. We apply our theory to the sets of $c$-convex functions on compact Hausdorff spaces, to the set of lower semicontinuous (convex) functions on a Hausdorff topological vector space and to 1-Lipschitz functions of complete metric spaces. The latter application is obtained using properties of the horoboundary of a metric space.
@@
Let $\mathscr{X}$ and $\mathscr{Y}$ be nonempty sets. Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a subset of $\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. $\overline{\mathbb{R}}^{\mathscr{Y}}$), with both $\mathscr{F}$ and $\mathcal{G}$ being proper and separating, stable by arbitrary suprema and addition of scalars. Let $J$ be a (max,+)-isomorphism from $\mathscr{F}$ onto $\mathcal{G}$ and define $e_{x}=\sup_{u\in\mathcal{G}}u(\cdot)-u(x)$ and $e^{\prime}_{y}=\sup_{v\in\mathscr{F}}v(\cdot)-v(y)$. Then the following statements are equivalent:
@@
The precise statement is given in ˜21, proper and separating corresponding to ˜1, and the functions $e_{x}$ being introduced in (18). The other properties of $g$ and $\phi$ depend obviously on the specific $\mathscr{F}$ and $\mathcal{G}$ studied, some characterizations obtained are collected in Table˜1. The particularity of our approach is that we encompass all these functions spaces within a single abstract formalism of sup-stable spaces. ˜21 establishing (1) is only a posteriori adapted to characterize the $g$ and $\phi$ for each specific space.
@@
1-Lipschitz functions constant isometry ˜33
@@
˜41
Concerning order reversing isomorphisms $T$ between sets $\mathscr{F}$ and $\mathcal{G}$ commuting with the addition of scalars, we characterize their existence through kernels $b$ in ˜10. Owing to the immediate ˜11, the uniqueness of $T$ up to basic transformations is instead a consequence of results like ˜1.
On $c$-convex functions over a compact set, we show in ˜27 that the existence of a (max,+)-isomorphism between $b$-convex and $c$-convex functions is equivalent to a “conjugacy” formula (25) satisfied by the kernels. Concerning 1-Lipschitz functions, the results rely on the compactification of metric spaces by horofunctions, and in particular on properties of Busemann functions (˜35). We also allow for weak distances, i.e. potentially assymetric and with negative values.
Similarly to the proof of the Banach-Stone theorem, see e.g. the lecture notes [GJ02], the key idea in our proofs is to advocate for inf and sup-irreducibility as a good notion of “extremality” (˜4) and to show that these “extremal” points are preserved by order isomorphisms (˜13) and generate the whole space, in a Krein–Milman fashion (˜19 and ˜25).
Beyond (max,+)-isomorphisms, we study the order isomorphisms of sets containing the indicator functions of points, such as the set of lower semicontinuous functions on Hausdorff topological spaces (˜29) and its subset of convex functions (˜41). In these cases, indicator functions play a major role in the characterization. We in particular obtain the extension of the theorems of [AAM09, IRS15] in the generality of locally convex Hausdorff spaces.
@@
Some of our results have precursors when $\mathscr{X}$ is finite, in the setting of tropical geometry – the (max,+) structure considered here being the main example of tropical algebraic structure. In particular, the notion of irreducibility was used in [BSS07, GK07] when studying the max-plus analogue of the representation of convex sets in terms of extreme points or extreme rays. Concerning infinite sets, the literature can be grouped according to the function spaces studied:
First, on convex functions, our focus meets that of the research program set by S. Artstein-Avidan and V. Milman on the characterization of some operations on subsets of convex functions. In their seminal paper [AAM09], they have shown that, for $\mathcal{G}$ the space of convex lower semicontinuous functions over $\mathbb{R}^{d}$, the only possible anti-involutions are compositions of the Fenchel transform with affine transformations. This result was further extended to Banach spaces by [IRS15], followed by [CL21] who characterized for which Banach spaces the Fenchel transform is an anti-involution. Using techniques specific to the convex case, a further generalization to convex subsets of locally convex Hausdorff spaces was obtained in [Leu23, Theorem 2.13], which contains our ˜41. We refer to [IRS15, AASW23] for a list of other developments on $s$-concave functions or convex bodies. Underlying all of these works on convex functions is a fundamental result of affine geometry that, in dimensions larger than 2, transformations preserving straight lines are affine. The notable difference of our approach is that we consider spaces beyond convex functions on vector spaces, so in general we do not have linear structures.
@@
Finally, an important application of ours is to the set of Lipschitz functions on a complete metric space with Lipschitz constant at most 1. Most of the interest for spaces of Lipschitz functions has been for their Banach structure, but they also have a rich lattice structure as advocated by [Wea94]. This induces quite some polysemy on the notion of isomorphism. For instance [CCD19] studies Banach space isomorphisms between spaces of Lipschitz functions. Weaver leveraged instead the inf/sup-lattice properties of the vector spaces of Lipschitz functions to characterize linear Banach space isomorphisms preserving Riesz-norms [Wea94]. Banach–Stone-like theorems for Lipschitz functions have also been shown for linear isomorphisms such as [JVCVV09, Theorem 4.1] and nonlinear [CSCS11, Theorem 1] and references within. Finally Leung et al. for linear, in [LL13], and nonlinear order isomorphisms, in [LT16], designed an abstract framework to describe order isomorphisms on “near vector lattices”. However all the aforementioned authors rely extensively on vector subspaces of the space of continuous functions, while we consider sup-stable sets of functions. The closest research to our work was done by [CSJ17, DJVM20] for $1$-Lipschitz functions on metric spaces with non-symmetric distance with the assumption that the isomorphisms also preserve convex combinations. Nevertheless, none of the authors uses the notions of inf or sup-irreducible functions. Moreover, a central idea in our proof is the use of the compactification of a metric space by horofunctions. We exploit the fact that Busemann points (limits of geodesics) are precisely the non-trivial sup-irreducible elements of the space of $1$-Lipschitz functions [AGW09].
There is a separate direction of research to characterize cones for which every order isomorphism is affine (in $f$). We refer to [LvGvI20, Wal19] for some technical conditions ensuring this. They unfortunately are not sufficient on spaces like that of continuous functions if not paired with homogeneity assumptions on $J$, in line with the work of [Sch79].
@@
In Section˜1 we introduce the notions of irreducibility and of isomorphisms that we consider. Section˜2 pertains to our main results on order isomorphisms, in particular when commuting with constants, with Section˜2.3 devoted to $c$-convex functions. In Section˜3, we detail the consequences of these results for continuous and lower semicontinuous functions (Section˜3.1), for 1-Lipschitz functions (Section˜3.2), and finally for convex lower semicontinuous functions (Section˜3.3).
@@ 1 Preliminaries
Notation: The extended real line is denoted by $\overline{\mathbb{R}}=[-\infty,+\infty]$, we also use the notations $\mathbb{R}_{\top}=(-\infty,+\infty]$, $\mathbb{R}_{\bot}=[-\infty,+\infty)$ and $\mathbb{R}_{>0}=(0,+\infty)$. Given any sets $\mathscr{X}$ and $E$, we denote by $E^{\mathscr{X}}$ the set of functions $f$ from $\mathscr{X}$ to $E$ and $\operatorname*{Id}_{\mathscr{X}}:\mathscr{X}\to\mathscr{X}$ the identity map. When $\mathscr{X}$ is equipped with a topology, a function $f:\mathscr{X}\rightarrow\overline{\mathbb{R}}$ is said to be lower semicontinuous (l.s.c.) if its epigraph is a closed subset of $\mathscr{X}\times\overline{\mathbb{R}}$. Its domain is defined by $\operatorname*{Dom}(f)=\{x\in\mathscr{X}\,|\,f(x)<+\infty\}$. A function $f$ is said to be proper if its domain is non-empty and if it does not take the value -$\infty$. For $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ and some $x\in\mathscr{X}$, if there exists a function $g\in\mathcal{G}$ such that $g(x)\in\mathbb{R}$, we say that $\mathcal{G}$ is proper at $x$, and more simply that $\mathcal{G}$ is proper if it is proper at all $x\in\mathscr{X}$. When $\mathscr{X}$ and $\mathscr{Y}$ are topological vector spaces, the set $\mathscr{X}^{*}$ denotes the dual space of $\mathscr{X}$ of continuous linear forms from $\mathscr{X}$ to $\mathbb{R}$ and $\mathcal{L}(\mathscr{X},\mathscr{Y})$ refers to the space of continuous linear operators from $\mathscr{X}$ to $\mathscr{Y}$. We write the duality product over $\mathscr{X}^{*}\times\mathscr{X}$ as $\langle\cdot,\cdot\rangle:\mathscr{X}^{*}\times\mathscr{X}\to\mathbb{R}$.
@@
We say that a subset $\mathcal{G}$ of $\overline{\mathbb{R}}^{\mathscr{X}}$ is sup-stable (resp. finitely sup-stable) if it is stable under arbitrary sups (resp. pairwise sups), and that it is a complete subspace of $\overline{\mathbb{R}}^{\mathscr{X}}$ if it is stable under arbitrary sups and addition of scalars. For $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$, we define its sup and inf-closures as
$\overline{\mathcal{G}}^{\sup}$ $:=\{\sup_{\alpha\in\mathcal{A}}h_{\alpha}\,|\,\mathcal{A}\text{ an index set},\,\{h_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{G}\},$
$\overline{\mathcal{G}}^{\inf}$ $:=\{\inf_{\alpha\in\mathcal{A}}h_{\alpha}\,|\,\mathcal{A}\text{ an index set},\,\{h_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{G}\}.$
@@
Relative infima underline that infima or suprema can be done either w.r.t. the absolute (entry-wise) supremum of $\overline{\mathbb{R}}^{\mathscr{X}}$ or its relative counterpart in $\mathcal{G}$.
@@
Let $\mathcal{G}$ be a subset of $\overline{\mathbb{R}}^{\mathscr{X}}$. We say that a point $f\in\mathcal{G}$ is sup-irreducible if, for all $g,h\in\mathcal{G}$,
@@
[…]
@@
[…]
@@
[…]
@@
    a (max,+)-isomorphism if it is a max-isomorphism and if we have $J(f+\lambda)=Jf+\lambda$ for $\lambda\in\mathbb{R}$, assuming that $\mathcal{G}$ and $\mathscr{F}$ are complete subspaces of $\overline{\mathbb{R}}^{\mathscr{X}}$;
@@
    an anti-involution if $\mathcal{G}\rightarrow\mathcal{G}, JJ$$J=\operatorname*{Id}_{\mathcal{G}}$ and $J$ is order reversing.
@@
Let $\mathscr{F}\subset\overline{\mathbb{R}}^{\mathscr{Y}}$ and $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ be sup-stable sets. Then the order isomorphisms $J:\mathscr{F}\rightarrow\mathcal{G}$ commute with an arbitrary suprema, and thus coincide with max-isomorphisms. Moreover, if $J:\mathscr{F}\rightarrow\mathcal{G}$ is a max-isomorphism, then so is $J^{-1}:\mathcal{G}\rightarrow\mathscr{F}$. Anti-isomorphisms $\bar{F}:\mathscr{F}\to\mathcal{G}$ send relative infima of $\mathscr{F}$ onto suprema of $\mathcal{G}$, i.e. $\bar{F}(\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha})=\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha})$ for any $\{f_{\alpha}\}_{\alpha\in\mathscr{A}}\subset\mathscr{F}$.
@@
So $\bar{F}(\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha})=\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha})$.
@@
A map $B:\overline{\mathbb{R}}^{\mathscr{Y}}\rightarrow\overline{\mathbb{R}}^{\mathscr{X}}$ is said to be $\overline{\mathbb{R}}_{\max}$-sesquilinear if $B(\inf\{f_{i}\}_{i\in I})=\sup\{Bf_{i}\}_{i\in I}$ and $B(f+\lambda)=Bf-\lambda$ (with $\infty$ absorbing on the l.h.s.  i.e. $f+\infty=+\infty$, and -$\infty$ absorbing on the r.h.s., i.e. $Bf-\infty=$-$\infty$), for any finite index set $I$ and $\lambda\in\overline{\mathbb{R}}$; we say in addition that $B$ is continuous if $B(\inf\{f_{i}\}_{i\in I})=\sup\{Bf_{i}\}_{i\in I}$ holds even for infinite families.
@@
A map $\bar{B}:\overline{\mathbb{R}}^{\mathscr{Y}}\rightarrow\overline{\mathbb{R}}^{\mathscr{X}}$ is $\overline{\mathbb{R}}_{\max}$-sesquilinear and continuous if and only if there exists a kernel $b:\mathscr{X}\times\mathscr{Y}\rightarrow\overline{\mathbb{R}}$ such that $\bar{B}f(x)=\sup_{y\in\mathscr{Y}}b(x,y)-f(y)$ for all $f\in\overline{\mathbb{R}}^{\mathscr{Y}}$ (with -$\infty$ absorbing). Moreover in this case $b$ is uniquely determined by $\bar{B}$ as $b(\cdot,y)=\bar{B}\delta^{\top}_{y}$.
Hence the relation between $\overline{\mathbb{R}}_{\max}$-sesquilinear and continuous maps, a.k.a. (Fenchel-Moreau) conjugations, and kernels is one-to-one, so it is equivalent to study a map or its kernel. Moreover ˜9 allows to write
@@
where $b:\mathscr{X}\times\mathscr{Y}\rightarrow\overline{\mathbb{R}}$ is the kernel associated with $\bar{B}$. Notice also that any complete subspace $\mathcal{G}$ of $\overline{\mathbb{R}}^{\mathscr{X}}$ is a range for the specific choice $\mathscr{Y}=\mathcal{G}$ and $b(\cdot,g)=g(\cdot)$, since
$\mathcal{G}=\{\sup_{g\in\mathcal{G}}a_{g}+g(\cdot)\,|\,a_{g}\in\mathbb{R}_{\bot}\}.$
@@
with the convention that -$\infty$ is absorbing. Interesting examples of sets $\mathcal{G}=\operatorname*{Rg}(B)$ are that of lower semicontinuous $C$-semiconvex functions for $b(x,y)=-\frac{C}{2}\|x-y\|^{2}$ for some $C>0$, or of 1-Lipschitz functions for b(x,y)=d(x,y)b(x,y)=-d(x,y). We use a terminology rooted in abstract convexity theory, however these functions are nothing else than the $c$-convex functions in optimal transport theory, taking c(x,y)=b(x,y)c(x,y)=-b(x,y). The key relation is that $\bar{B}=\bar{B}\bar{B}^{\circ}\bar{B}$, see e.g.  [AGK05, p.3, Theorem 2.1, Example 2.8]. We provide now a characterization relating ranges and anti-isomorphisms.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a complete subspace of $\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. $\overline{\mathbb{R}}^{\mathscr{Y}}$). Then the following statements are equivalent: